OpenStudy (anonymous):

Use mathematical induction to prove that the formula is true for every positive integer n. 1. 2+4+6+8+....+(2n)=n(n+1) 2. 1+4+7+10+.....+(3n-2)=n(3n-1)/2 3.1^3+2^3+3^3+4^3.....+(n^3)= (n^2(n+1)^2)/4 4. 3+7+11+15+.....+(4n-1)= n(2n+1) 5. 1/(1*2*3)+1/(2*3*4)+ 1/(3*4*5)+1(4*5*6)+.....+(1/(n(n+1)(n+2))=n(n+3)/4(n+1)(n+2))

OpenStudy (anonymous):

Ok i would try the 3rd one for you so you will get the idea of it. so we should prove 1^3+2^3+3^3+4^3.....+(n^3)= (n^2(n+1)^2)/4 first we should check if its correct for n=1 L.H.S.=1 R.H.S.=1(2^2)/4=1 so its correct for n=1. assume it would be true when n=p so 1^3+2^3+3^3+4^3.....+(p^3)= (p^2(p+1)^2)/4 then lets try to show that its true for n=p+1 too like this 1^3+2^3+3^3+4^3.....+(p^3)= (p^2(p+1)^2)/4 adding (p+1)^3 for both sides 1^3+2^3+3^3+4^3.....+(p^3) +(p+1)^3= (p^2(p+1)^2)/4 +(p+1)^3 =(p+1)^2[p^2+4(p+1)]/4 =(p+1)^2[p^2+4p+4)]/4 =(p+1)^2[p+2^2)]/4 so it seems if we assume its correct for n=p it would be true to n=p+1 too. and we showed that its correct for n=1 in the beginning. so according to mathematical induction the result is true for every possible integer n * its like a chain of domions

OpenStudy (anonymous):

Is there any trick to reducing the equation to get it to look like the beginning one?

OpenStudy (anonymous):

I can figure everything but the last step is hardest for me

OpenStudy (anonymous):

Thnaks you mean reducing the equation? well i take common terms ( or what do u call them) out so the equation becomes simpler

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

you are welcome :)